# Neuroplastic Expansion in Deep Reinforcement Learning

### v1

@@ Abstract
The loss of plasticity in learning agents, analogous to the solidification of neural pathways in biological brains, significantly impedes learning and adaptation in reinforcement learning due to its non-stationary nature. To address this fundamental challenge, we propose a novel approach, Neuroplastic Expansion (NE), inspired by cortical expansion in cognitive science. NE maintains learnability and adaptability throughout the entire training process by dynamically growing the network from a smaller initial size to its full dimension. Our method is designed with three key components: (1) elastic neuron generation based on potential gradients, (2) dormant neuron pruning to optimize network expressivity, and (3) neuron consolidation via experience review to strike a balance in the plasticity-stability dilemma. Extensive experiments demonstrate that NE effectively mitigates plasticity loss and outperforms state-of-the-art methods across various tasks in MuJoCo and DeepMind Control Suite environments. NE enables more adaptive learning in complex, dynamic environments, which represents a crucial step towards transitioning deep reinforcement learning from static, one-time training paradigms to more flexible, continually adapting models.
@@ 1 Introduction
In the realm of neuroscience, it has been observed that biological agents often experience a diminishing ability to adapt over time, analogous to the gradual solidification of neural pathways in the brain (Livingston, 1966). This phenomenon, typically known as the loss of plasticity (Mateos-Aparicio and Rodríguez-Moreno, 2019), significantly affects an agent’s capacity to learn continually, especially when agents learn by trial and error in deep reinforcement learning (deep RL) due to the non-stationarity nature. The declining adaptability throughout the learning process can severely hinder the agent’s ability to effectively learn and respond to complex or non-stationary scenarios (Abbas et al., 2023). This limitation presents a fundamental obstacle to achieving sustained learning and adaptability in artificial agents, which echoes the plasticity-stability dilemma (Abraham and Robins, 2005) observed in biological neural networks.
@@
There have been several recent studies highlighting a significant loss of plasticity in deep RL (Kumar et al., 2020, Lyle et al., 2022), which substantially restricts the agent’s ability to learn from subsequent experiences (Lyle et al., 2023, Ma et al., 2023). The identification of primacy bias (Nikishin et al., 2022) further illustrates how agents may become overfitted to early experiences, which inhibits learning from subsequent new data. The consequences of plasticity loss further impede deep RL in continual learning scenarios, where the agent struggles to sequentially learn across a series of different tasks (Dohare et al., 2024).
@@
Research on addressing plasticity loss in deep RL is still in its early stages, with recent approaches including parameter resetting (Nikishin et al., 2022, Sokar et al., 2023) (or its advancement with random head copies (Nikishin et al., 2024)), and several implementation-level techniques like normalization, activation functions, weight clipping, and batch size adjustments (Obando Ceron et al., 2023, Nauman et al., 2024, Elsayed et al., 2024). However, reset-based methods often lead to performance instability and training inefficiency due to the need for period re-training, while implementation-level modifications lack generalizability and do not directly target the plasticity issue. The field currently lacks a unified methodology that can effectively address plasticity loss while maintaining training stability and efficiency across varying environments.
@@
Humans adapt to environmental changes and novel experiences through cortical cortex expansion in cognitive science (Hill et al., 2010). This process involves the gradual activation of additional neurons and the formation of new connections to facilitate the ability to learn continually. Drawing inspiration from this biological mechanism, we propose a novel perspective – Neuroplastic Expansion, which can help maintain plasticity in deep RL. The key insight is that an agent starting learning with a smaller network and dynamically growing to a larger size (ultimately reaching the original static network dimension) can effectively tackle plasticity loss by maintaining a high level of elastic neurons throughout training (Figure 1). We first provide empirical evidence in (§4.1) validating its potential for mitigating plasticity loss and improving final performance in certain cases, even with a naive incremental expansion.
@@
Building upon this insight, we systematically introduce Neuroplastic Expansion (NE), a simple yet effective mechanism to maximize the benefits of incremental growth training (§4.2). (NE) adds high-quality elastic candidates based on potential gradients (Evci et al., 2020). However, this will lead to increased computation cost for always enlarging the network, and cannot fully leverage its expressivity as neurons may turn dormant during the course of training. To address this challenge, (NE) introduces a reactivation process for dormant neurons, pruning them and potentially reintroducing them as candidates in subsequent growth stages, thereby better utilizing the network expressivity and enhancing the policy’s sustainable learning ability. Unlike previous reset-based approaches (either rest last layers or introducing copies of heads or reinitializing some parts in the network) that risk forgetting and require periodic re-training, NE maintains learning continuity in a smoother way. To further mitigate potential instability from continuous topology adjustments, we introduce an effective Experience Review technique for consolidating neurons by adaptively reviewing prior knowledge during the later training stages to uphold policy stability, striking a balance in the stability-plasticity dilemma. Extensive experiments across standard and continual RL tasks demonstrate its effectiveness in various scenarios, showcasing its ability to maintain plasticity while ensuring policy stability.
@@
  • We introduce a novel mechanism, Neuroplastic Expansion (NE), to mitigate the plasticity loss in deep RL.
  • We develop effective elastic neuron generation and dormant neuron pruning mechanism for better network capacity utilization, and a neuron consolidation technique for preventing forgetting helpful reusable knowledge.
@@ 2 Background
The reinforcement learning problem can be typically formulated by a Markov decision process (MDP) represented as a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$, with $\mathcal{S}$ denoting the state space, $\mathcal{A}$ the action space, $\mathcal{P}$ the transition dynamics: $\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$, $\mathcal{R}$ the reward function: $\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, and $\gamma\in[0,1)$ the discount factor. The agent interacts with the unknown environment with its policy $\pi$, which is a mapping from states to actions, and aims to learn an optimal policy that maximizes the expected discounted long-term reward. The state-action value of $s$ and $a$ under policy $\pi$ is defined as $Q^{\pi}(s,a)=\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(s_{t},a_{t})|s_{0}=s,a_{0}=a]$. In actor-critic methods, the actor $\pi_{\phi}$ and the critic $Q_{\theta}$ are represented using neural networks as function approximators with parameters $\phi$ and $\theta$ (Fujimoto et al., 2018, Haarnoja et al., 2018). The critic network is updated by minimizing the temporal difference loss, i.e., $\mathcal{L}_{Q}(\theta)=\mathbb{E}_{D}\big{[}(Q_{\theta}(s,a)-Q^{\mathcal{T}}_{\theta}(s,a))^{2}\big{]}$, where $Q^{\mathcal{T}}(s,a)$ denotes the bootstrapping target $\mathcal{R}(s,a)+\gamma Q_{\bar{\theta}}(s^{\prime},\pi_{\bar{\phi}}(s^{\prime}))$ computed using target network parameterized by $\bar{\phi}$ and $\bar{\theta}$  based on data sampled from a replay buffer $D$. The actor network $\phi$ is typically updated to maximize the Q-function approximation according to $\nabla_{\phi}J(\phi)=\mathbb{E}_{D}\left[\left.\nabla_{a}Q_{\theta}(s,a)\right|_{a=\pi_{\phi}(s)}\nabla_{\phi}\pi_{\phi}(s)\right]$.
Dormant Neuron & Activated Neuron
The concepts of dormant and activated neurons are crucial in assessing network plasticity. Dormant neurons are defined as those whose output falls below a certain threshold $\tau$ in the neural network (Xu et al., 2023), providing a quantitative measure of network plasticity. Formally, a neuron $i$ in layer $l$ is considered $\tau$-dormant when $f(l_{i})\leq\tau$ (Sokar et al., 2023), where
$f(l_{i})=\frac{\mathbb{E}_{x\in Id}|h^{l}_{i}(x)|}{\frac{1}{H^{l}}\sum_{k\in h}\mathbb{E}_{x\in Id}|h^{l}_{k}(x)|},$ (1)
with $Id$ denoting the input distribution. The activated neuron ratio, defined as the complement of the dormant neuron ratio (i.e., $1-$ dormant neuron ratio), serves as an alternative metric for plasticity. Activated neuron ratio is widely used in plasticity visualization (Ma et al., 2023) due to its intuitive interpretation (higher values correspond to better performance).
3 Related Work
Plasticity loss in RL
Recent studies have found that agents guided by the non-stationary objective characteristic of RL suffer catastrophic degradation (Lyle et al., 2023) and ultimately reduce the ability to perform continual learning. This phenomenon is known as plasticity loss. Behaviorally, the loss of plasticity during training can manifest as, for instance, an inability to obtain an effective gradient to guide policy adaptation in fine-tuning (Abbas et al., 2023), limiting potential transfer performance. While the research on this topic is still in its primary stages, there are several techniques that have been demonstrated to mitigate the loss of plasticity in deep RL. Lyle et al. (2024) empirically find that network normalization particularly layer regularization (LN), is very effective in mitigating plasticity loss. Recently, Nikishin et al. (2024) proposed a plug-in module that can be used to improve continual learning ability and as a metric of plasticity evaluation, and introduced copies of random heads for injecting plasticity. In pixel-based deep RL, data augmentation (Ma et al., 2023) and batch size reduction (Obando Ceron et al., 2023) have been analyzed in reducing plasticity loss.
As highlighted by (Abbas et al., 2023), plasticity loss diminishes a network’s ability to effectively minimize training loss on new data (trainability) and reduces its capacity to generalize to unseen data (generalizability). In the former case, Nikishin et al. (2022) showed that deep RL agents overfit early experiences, which limits their ability to learn with new experiences later on. Consequently, the agent performs suboptimally in long-term learning. This effect is called the primacy bias. Recent studies have found that techniques focusing on mitigating primacy bias such as parameter resetting (Schwarzer et al., 2023, Kim et al., 2024), mitigating gradient starvation (Dohare et al., 2024), and constraining deviations from initial weights (Kumar et al., 2023, Lewandowski et al., 2024) can also be effective in mitigating plasticity loss. Notably, the resetting strategy has been shown to work at scale, with applications in LLM fine-tuning (Noukhovitch et al., 2024). The reason why the two mentioned domains can share a solution is likely that there is a common incentive, that is, the ability of network parameters to learn cannot be maintained by only following the traditional training processes. Thus, this paper proposes a new general framework from the perspective of training mechanisms to simultaneously alleviate the primacy bias and plasticity loss issues in deep RL with a better trade-off between stability and plasticity.
Pruning in RL
Sokar et al. (2021) showed that training the deep RL policy with a changing topology is difficult due to training instability. Since then, this challenging topic has been well-studied. Policy Pruning and Shrinking (PoPs) (Livne and Cohen, 2020) obtain a sparse deep RL agent with iterative parameter pruning. Graesser et al. (2022), Tan et al. (2022) attempt to train a sparse neural network from scratch without pre-training a dense teacher. Existing methods mainly focus on obtaining a small topology to improve training efficiency. However, our paper explores whether growing the topology from small to large can mitigate the plasticity loss in RL agents rather than training with the maximum number of parameters. In deep RL, the model size commonly utilized is typically small and may encounter policy collapse after scaling up (Schwarzer et al., 2023, Ceron et al., 2024a). Thus, sparse training techniques that have proven beneficial for significantly reducing training costs in fully supervised domains may not yield similar advantages in deep RL in terms of computation cost reduction. Instead, we provide a new perspective and focus on making deep RL agents better use of the growing network size to further improve their learning ability.
@@ 4 Neuroplastic Expansion RL
In this section, we begin by discussing the insights of Neuroplastic Expansion RL empirically, and analyze its effect on mitigating loss of plasticity in Section 4.1. Then, we systematically present details of our method in Section 4.2.
@@ 4.1 Illustration of Insights of Dynamic Growing RL
There have been several recent studies investigating loss of plasticity in supervised learning (Lyle et al., 2023, Lewandowski et al., 2024), which corresponds to the phenomenon where neural networks gradually lose the ability to adapt and learn from new experiences (Lin et al., 2022). This is further exacerbated in RL (Sokar et al., 2023, Nikishin et al., 2024), due to the non-stationary nature and the tendency to overfit prior knowledge, resulting in suboptimal policies. A predominant approach to tackle plasticity loss is based on resetting (Nikishin et al., 2022), where the agent resets the last few layers of its neural network periodically throughout training. However, although effective, this kind of approach typically forgets helpful and reusable knowledge which is critical for learning, and thus the agent experiences a drop in performance once resetting is applied.
@@
One of the major causes for agents to adapt to novel circumstances is based on the growth of the human cortex (Hill et al., 2010). Motivated by this insight, we propose a novel perspective that RL agents can maintain high plasticity through dynamic neural expansion mechanisms, sustaining the policy’s continual adaptation ability from new experiences. Initially, data collected by a random policy may be of lower quality and therefore require less network expressivity (Burda et al., 2018, Zhelo et al., 2018). As the agent improves, it tends to need a more expressive network for fitting complex value estimates and policies. Maintaining neuronal vitality and regeneration during training can significantly mitigate plasticity loss, as it allows for ongoing adaptation to new information.
@@
Figure 2: Comparison of normalized performance (left) and the ratio of active neurons (right) for vanilla TD3 (Fujimoto et al., 2018) and its variant with a naive random growth network and our Neuroplastic Expansion method, across varying network capacities. Experiments are conducted with 7 independent seeds.
@@
Motivated by this insight, we propose a novel angle considering Neuroplastic Expansion, where the agent starts with a smaller network that evolves to a larger one throughout the learning process. We first demonstrate a naive implementation of this idea and will systematically discuss our methodology in Section 4.2. Concretely, we consider a small-to-large neural expansion approach, where the capacity of initial networks is $20\%$ of the size of a typical full TD3 (Fujimoto et al., 2018) network (i.e., uniformly selected $20\%$ of neurons from each layer in the full TD3). To validate the potential of this novel perspective, we first consider the simplest manner for growing the network by uniformly adding $k$ new connections in the current critic topology $\breve{\theta}\in\theta$ and current actor topology $\breve{\phi}\in\phi$ for each layer $l\in N$ and participate in gradient propagation according to $\mathbb{I}_{grow}^{l}=\text{Random}_{i\notin\breve{\theta}_{l}}(\theta_{l},k)+\text{Random}_{i\notin\breve{\phi}_{l}}(\phi_{l},k)$.
@@
We build this idea upon TD3 and conduct fair comparisons with vanilla TD3 (i.e., with static actor and critic networks) with different capacities. We compared the performance of the dynamic growing network with a naive topology growth strategy, with the typical ones using static networks based on TD3 (Fujimoto et al., 2018) in the HalfCheetah environment. To remove the difference in fitting ability caused by model capacity and capture the impact of gradual topology growth, the number of termination parameters of the topology growth is set to be consistent with the traditional static parameter methods. Therefore, the comparison with the vanilla TD3 algorithm is fair given the same network capacity at the end of training.
@@
From the left part in Figure 2, we observe that even a naive incremental network growth can bring relatively noticeable gains to the agent considering larger network capacity. The right part in Figure 2 demonstrates the ratio of active neurons (Sokar et al., 2023) for each algorithm during training. This metric is used to evaluate RL agents regarding whether they maintain expressive ability. As shown, topology growth can effectively alleviate neuron deactivation and thus maintain the ability of policy learning to mitigate the loss of plasticity and alleviate the primacy bias.

### v3

@@ Abstract
The loss of plasticity in learning agents, analogous to the solidification of neural pathways in biological brains, significantly impedes learning and adaptation in reinforcement learning due to its non-stationary nature. To address this fundamental challenge, we propose a novel approach, Neuroplastic Expansion (NE), inspired by cortical expansion in cognitive science. NE maintains learnability and adaptability throughout the entire training process by dynamically growing the network from a smaller initial size to its full dimension. Our method is designed with three key components: (1) elastic topology generation based on potential gradients, (2) dormant neuron pruning to optimize network expressivity, and (3) neuron consolidation via experience review to strike a balance in the plasticity-stability dilemma. Extensive experiments demonstrate that NE effectively mitigates plasticity loss and outperforms state-of-the-art methods across various tasks in MuJoCo and DeepMind Control Suite environments. NE enables more adaptive learning in complex, dynamic environments, which represents a crucial step towards transitioning deep reinforcement learning from static, one-time training paradigms to more flexible, continually adapting models. We make our code publicly available.
@@ 1 Introduction
In the realm of neuroscience, it has been observed that biological agents often experience a diminishing ability to adapt over time, analogous to the gradual solidification of neural pathways in the brain (Livingston, 1966). This phenomenon, typically known as the loss of plasticity (Mateos-Aparicio & Rodríguez-Moreno, 2019), significantly affects an agent’s capacity to learn continually, especially when agents learn by trial and error in deep reinforcement learning (deep RL) due to the non-stationarity nature. The declining adaptability throughout the learning process can severely hinder the agent’s ability to effectively learn and respond to complex or non-stationary scenarios (Abbas et al., 2023). This limitation presents a fundamental obstacle to achieving sustained learning and adaptability in artificial agents, which echoes the plasticity-stability dilemma (Abraham & Robins, 2005) observed in biological neural networks.
@@
There have been several recent studies highlighting a significant loss of plasticity in deep RL (Kumar et al., 2020; Lyle et al., 2022), which substantially restricts the agent’s ability to learn from subsequent experiences (Lyle et al., 2023; Ma et al., 2023). The identification of primacy bias (Nikishin et al., 2022) further illustrates how agents may become overfitted to early experiences, which inhibits learning from subsequent new data. The consequences of plasticity loss further impede deep RL in continual learning scenarios, where the agent struggles to sequentially learn across a series of different tasks (Dohare et al., 2024).
@@
Research on addressing plasticity loss in deep RL is still in its early stages, with recent approaches including parameter resetting (Nikishin et al., 2022; Sokar et al., 2023) (or its advancement with random head copies (Nikishin et al., 2024)), and several implementation-level techniques like normalization, activation functions, weight clipping, and batch size adjustments (Obando Ceron et al., 2023; Nauman et al., 2024; Elsayed et al., 2024). However, reset-based methods often lead to performance instability and training inefficiency due to the need for period re-training, while implementation-level modifications lack generalizability and do not directly target the plasticity issue. The field currently lacks a unified methodology that can effectively address plasticity loss while maintaining training stability and efficiency across varying environments.
@@
Humans adapt to environmental changes and novel experiences through cortical cortex expansion in cognitive science (Welker, 1990; Hill et al., 2010). This process involves the gradual activation of additional neurons and the formation of new connections to facilitate the ability to learn continually. Drawing inspiration from this biological mechanism, we propose a novel perspective – Neuroplastic Expansion, which can help maintain plasticity in deep RL. The key insight is that an agent starting learning with a smaller network and dynamically growing to a larger size (ultimately reaching the original static network dimension) can effectively tackle plasticity loss by maintaining a high level of elastic neurons throughout training (Figure 1). Elastic neurons means neurons that have plasticity. In this paper, the words elastic and plastic are interchangeably. We first provide empirical evidence in (§3.1) validating its potential for mitigating plasticity loss and improving final performance in certain cases, even with a naive incremental expansion. We provide a detailed description of plasticity loss and pruning for RL in Appendix A.
@@
Building upon this insight, we systematically introduce Neuroplastic Expansion (NE), a simple yet effective mechanism to maximize the benefits of incremental growth training (§3.2). NE adds high-quality elastic candidates based on potential gradients (Evci et al., 2020). However, this will lead to increased computation costs for always enlarging the network, and cannot fully leverage its expressivity as elastic neurons, which are activated and can be updated to fit the new data, may turn dormant during the course of training. As visualized in Appendix F.6, dormant neurons can diminish representational capacity and lead to suboptimal behavior.
To address this challenge, NE introduces a reactivation process for dormant neurons, pruning them and potentially reintroducing them as candidates in subsequent growth stages; thereby better utilizing the network expressivity and enhancing the policy’s sustainable learning ability.
Unlike previous reset-based approaches (such as resetting the last layers, introducing copies of heads or reinitializing parts of the network) that risk forgetting and require periodic re-training, NE maintains learning continuity in a smoother way. To further mitigate potential instability from continuous topology adjustments, we introduce an effective Experience Review technique that consolidates neurons by adaptively reviewing prior knowledge during the later training stages, ensuring policy stability and balancing the stability-plasticity dilemma. Extensive experiments on long-term training, continual adaptation, and vision RL tasks demonstrate the effectiveness of our method across various scenarios, showcasing its ability to maintain plasticity while ensuring policy stability.
@@
  • We introduce a novel mechanism, Neuroplastic Expansion (NE), to mitigate the loss of plasticity in deep RL.
  • We develop effective activated neuron generation and dormant neuron pruning mechanism for better network capacity utilization, and a neuron consolidation technique for preventing forgetting helpful reusable knowledge.
@@ 2 Background
The reinforcement learning problem can be typically formulated by a Markov decision process (MDP) represented as a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$, with $\mathcal{S}$ denoting the state space, $\mathcal{A}$ the action space, $\mathcal{P}$ the transition dynamics: $\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$, $\mathcal{R}$ the reward function: $\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, and $\gamma\in[0,1)$ the discount factor. The agent interacts with the unknown environment with its policy $\pi$, which is a mapping from states to actions, and aims to learn an optimal policy that maximizes the expected discounted long-term reward. The state-action value of $s$ and $a$ under policy $\pi$ is defined as $Q^{\pi}(s,a)=\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(s_{t},a_{t})|s_{0}=s,a_{0}=a]$.
In actor-critic methods, the actor $\pi_{\phi}$ and the critic $Q_{\theta}$ are represented using neural networks as function approximators with parameters $\phi$ and $\theta$ (Fujimoto et al., 2018; Haarnoja et al., 2018). The critic network is updated by minimizing the temporal difference loss, i.e., $\mathcal{L}_{Q}(\theta)=\mathbb{E}_{D}\big{[}(Q_{\theta}(s,a)-Q^{\mathcal{T}}_{\theta}(s,a))^{2}\big{]}$, where $Q^{\mathcal{T}}(s,a)$ denotes the bootstrapping target $\mathcal{R}(s,a)+\gamma Q_{\bar{\theta}}(s^{\prime},\pi_{\bar{\phi}}(s^{\prime}))$ computed using target network parameterized by $\bar{\phi}$ and $\bar{\theta}$  based on data sampled from a replay buffer $D$. The actor network $\phi$ is typically updated to maximize the Q-function approximation according to $\nabla_{\phi}J(\phi)=\mathbb{E}_{D}\left[\left.\nabla_{a}Q_{\theta}(s,a)\right|_{a=\pi_{\phi}(s)}\nabla_{\phi}\pi_{\phi}(s)\right]$.
Activated Neuron Ratio
Activated neurons are easily updated by new data. Therefore, the proportion of activated neurons in the network, i.e. activated neuron ratio is often considered positively correlated with plasticity Ma et al. (2023). It is defined as the proportion of neurons whose output exceeds a certain threshold $\tau$ in the neural network, effectively counting all neurons except dormant ones (Xu et al., 2023). Formally, a neuron $i$ in layer $l$ is considered activated when output $h_{i}(x)>0$ (in contrast to dormant neurons (Sokar et al., 2023)):
$f(l_{i})=\frac{\sum_{i\in l,x\in Id}\mathrm{1}(h_{i}(x)>0)}{N},$ (1)
where $Id$ denotes the input distribution. The relationship between the Activated Neuron Ratio curve and plasticity is as follows: The primary cause behind an agent’s plasticity loss is the rapid dormancy of numerous neurons as training progresses, which diminishes the model’s representational capacity (Sokar et al., 2023; Qin et al., ). Consequently, delaying the reduction in activated neurons is commonly considered as a way to mitigate plasticity loss, i.e, slowing the downward trend of the curve. The Activated Neuron Ratio is widely used for plasticity visualization (Ma et al., 2023) due to its intuitive interpretability.
@@ 3 Neuroplastic Expansion RL
In this section, we begin by discussing the insights of Neuroplastic Expansion RL empirically, and analyze its effect on mitigating the loss of plasticity in Section 3.1. Then, we systematically present the details of our method in Section 3.2.
@@ 3.1 Illustration of Insights of Dynamic Growing RL
There have been several recent studies investigating loss of plasticity in supervised learning (Lyle et al., 2023; Lewandowski et al., 2024), which corresponds to the phenomenon where neural networks gradually lose the ability to adapt and learn from new experiences (Lin et al., 2022). This is further exacerbated in RL (Sokar et al., 2023; Nikishin et al., 2024), due to the non-stationary nature and the tendency to overfit prior knowledge, resulting in suboptimal policies. A predominant approach to tackle plasticity loss is based on resetting (Nikishin et al., 2022), where the agent resets the last few layers of its neural network periodically throughout training. However, although effective, this kind of approach typically forgets helpful and reusable knowledge which is critical for learning, and thus the agent experiences a drop in performance once resetting is applied.
@@
A major cause for agents to adapt to novel circumstances is the growth of the human cortex (Hill et al., 2010). Motivated by this insight, we propose a novel perspective that RL agents can maintain high plasticity through dynamic neural expansion mechanisms, sustaining the policy’s continual adaptation ability based on new experiences. Initially, data collected by a random policy may be of lower quality and therefore require less network expressivity to fit (Burda et al., 2018; Zhelo et al., 2018). As the agent improves, it tends to require a more expressive network for fitting the more complex value estimates and policies. Maintaining neuronal vitality and regeneration during training can significantly mitigate plasticity loss, as it allows for ongoing adaptation to new information.
@@
Figure 2: Comparison of normalized performance (left) and the ratio of active neurons (right) for vanilla TD3 (Fujimoto et al., 2018) and its variants: with a naive random growth network and our NE method, across varying network capacities.
Motivated by this insight, we propose a novel angle, Neuroplastic Expansion, where the agent starts with a smaller network that gradually evolves into a larger one throughout the learning process. We first present a naive implementation of this idea and then systematically discuss our methodology in Section 3.2.
@@
Concretely, we adopt a small-to-large neural expansion approach, where the initial network capacity is $20\%$ of the size of a typical full TD3 (Fujimoto et al., 2018) network (i.e., uniformly selected $20\%$ of neurons from each layer of the full TD3 architecture). To validate the potential of this novel perspective, we first consider the simplest approach for growing the network by uniformly adding $k$ new connections to the current critic topology $\breve{\theta}\subset\theta$ and the current actor topology $\breve{\phi}\subset\phi$ for each layer $l\in N$, which then participate in gradient propagation according to $\mathbb{I}_{grow}=\text{Random}_{i\notin\breve{\theta}_{l}}(\theta_{l},k)\cup\text{Random}_{i\notin\breve{\phi}_{l}}(\phi_{l},k)$.
@@
We build this idea upon TD3 and conduct fair comparisons with vanilla TD3 (i.e., with static actor and critic networks) using different network capacities. We compare the performance of our dynamically growing network against a naive topology growth strategy and static networks in standard TD3 (Fujimoto et al., 2018) using the HalfCheetah environment. To ensure fairness in comparison, the final network size of NE is set to be identical to that of the other methods. Therefore, the comparison with the vanilla TD3 algorithm is fair given the same network capacity at the end of training.
@@
From the left side of Figure 2, we observe that even a naive incremental network growth strategy leads to noticeable performance improvements, particularly as network capacity increases. The right side of Figure 2 demonstrates the ratio of active neurons (Sokar et al., 2023) for each algorithm during training. This metric evaluates whether RL agents maintain their expressive capacity, which is positively correlated with plasticity retention. As shown, topology growth effectively reduces neuron deactivation, thereby preserving the agent’s ability to learn policies, mitigating plasticity loss, and alleviating primacy bias.